OpenStudy (ayyookyndall):

PLEASE HELP!!! Find the geometric mean of 5 and 8. Write the exact answer--- do not round.

OpenStudy (ayyookyndall):

@ganeshie8 @Kainui

OpenStudy (anonymous):

Multiply the two numbers and take the square root

OpenStudy (ribhu):

the geometric mean of two numbers a and b is squareroot of product of a&b i.e. \[\sqrt{ab}\] so in this case it would be \[\sqrt{5*8}\] = 6.32

OpenStudy (ribhu):

@ayyookyndall please see the solution

OpenStudy (ayyookyndall):

Oh okay. Can you guys help me with another question? @ribhu @wowok

OpenStudy (ribhu):

yeah which question?

OpenStudy (anonymous):

Sure! :)

OpenStudy (anonymous):

You ned to set up a ratio first \[x \div 4.3 = 14.8 \div 17.8\]

OpenStudy (anonymous):

After that, cross multiply.

OpenStudy (ayyookyndall):

Is it a?

OpenStudy (anonymous):

Ah I miss-typed the equation. It should be \[x \div 8.3 = 14.8 \div 17.8\]

OpenStudy (ribhu):

how? @wowok

OpenStudy (ayyookyndall):

Is my answer correct?

OpenStudy (ribhu):

yeah yours is correct @ayyookyndall

OpenStudy (anonymous):

Yes! Sorry, I totally messed up the equation the first time. It's nearly 1am where I live. Good job!

OpenStudy (ribhu):

the triangles would be similar hence 14.8/7.8 would be equal to x/8.3

OpenStudy (ayyookyndall):

Help with more?

OpenStudy (anonymous):

Go for it! :)

OpenStudy (ayyookyndall):

The measures of the angles of a triangle are in the extended ratio 2:2:8. What is the measure of the smallest angle? @wowok @ribhu

OpenStudy (anonymous):

Is there a diagram?

OpenStudy (ribhu):

let the angles be 2x, 2x and 8x and since the sum of all the angles in a triangle is 180 degrees so 2x+2x+8x = 180 =.12x=180 => x=15

OpenStudy (ayyookyndall):

No

OpenStudy (ribhu):

all the angles would be 2x = 30, 2x = 30, 8x = 120

OpenStudy (anonymous):

There is 12 parts then. So it has to equal 180

OpenStudy (anonymous):

so each 2 will be 30 and the 8 is 120

OpenStudy (ribhu):

yeah x would be any parameter which would make the angles in the ratio of 2:2:8

OpenStudy (ribhu):

@ayyookyndall did u get it

OpenStudy (ayyookyndall):

Okay. yes... A couple more?

OpenStudy (ayyookyndall):

Sorry its looks bad

OpenStudy (ribhu):

i am afraid the question is not clear @ayyookyndall

OpenStudy (ayyookyndall):

@ribhu

OpenStudy (ribhu):

yeah i wil try

OpenStudy (ribhu):

here triange APQ similar to triangle ABC

OpenStudy (ayyookyndall):

Here how about other one... You dont have to do it if you dont know it.

OpenStudy (ribhu):

AP/AB = AQ/AC => AP = 8, AB = 8+x, AQ = 12, AC = 12+18 so substituting the values we get 8/(8+x) = 12/30 so the equations becomes

OpenStudy (ayyookyndall):

Are you sure?

OpenStudy (ribhu):

240 = 96 + 12x (upon cross multiplication) x = 12 (final answer) @ayyookyndall

OpenStudy (ribhu):

yeah i am damn sure for this.